Tessie is twice as old as Norma. Three years ago, she was three times as old. How old are they now?
Answer by Maths68(1474)

Let
Present age of Tessie = t
Present age of Norma = n
Given
Tessie is twice as old as Norma
t=2n ..........(1)
Three years ago, means we have to subtract 3 from their present ages
t-3=n-3
then
Tessie was three times older than Norma
t-3=3(n-3)
t-3=3n-9
t=3n-9+3
t=3n-6 ............(2)
Put the value of t from (1) to (2)
t=3n-6 ............(2)
2n=3n-6
2n-3n=-6
-n=-6
Cancel out (-1) from both sides
n=6
Put the value of n in (1)
t=2n
t=2(6)
t=12

Present age of Tessie = t = 12 years
Present age of Norma = n = 6 years